5-An the economy of Pizzania (from Problem 4, bread lt bread - and cheese produced are sold both to the pizza com- pany for inputs in the production of pizzas and to consumers as final goods. The accompanying table summarizes the activities of the three companies Bread Cheese company company Pizza company Cost of inputs $0 $%50 (bread) $0 35 (cheese) Wages Value of output 25 30 75 100 60 200 a. Calculate GDP as the Value addedin production b. Calculate GDP as spending on final goods and services. c. Calculate GDP as factof Yncome
Answer #1

Answer 5

(a)

GDP using Value added method is given by:

GDP = Value of Output - Intermediate consumption

Here, Value of Output = Sum of Value of Output of all Company = 100 + 60 + 200 = 360

Intermediate Consumption = Total Cost of Inputs = 0 + 0 + 50 + 35 = 85

Hence, GDP = 360 - 85 = 275

(b)

Total Spending on final Good of Bread = 100 - 50 = 50. Because 50 of the bread is not Final

Total Spending on final Good of Cheese = 60 - 35 = 25. Because 35 of the Cheese is not Final

Total Spending on final Good(Pizza) = 200

Hence, GDP = 50 + 25 + 200 = 275

(c)

Finally Lets Calculate GDP using Factor Income

Factor income includes wages, Profits, etc.

Here wages = 25 + 30 + 75 = 130

Here Profit of Bread Company = 100 - 25 = 75

Profit of Cheese Company = 60 - 30 = 30

Profit of Burger Company = 200 - 75 - 50 - 35 = 40

Total GDP = 130 + 30 + 75 + 40 = 275